OpenStudy (anonymous):

Use the quadratic formula to solve -4x2 + 4x + 80 = 0 pls help

OpenStudy (mertsj):

First divide everything by -4

OpenStudy (mertsj):

x^2-x-20=0

OpenStudy (mertsj):

a=1, b=-1, c=-20

OpenStudy (mertsj):

\[x=\frac{1\pm \sqrt{1-4(-20)}}{2}\]

OpenStudy (mertsj):

Simplify. You will get 5 and -4

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

-4x^2 + 4X + 80 =0 is in the form of Ax^2 + Bx^2+C=0 from the quadratic eq, put a, b and c respectively. Don't forget, we have two x values here, with + and - signs.

OpenStudy (anonymous):

ok thank for the help